OpenStudy (anonymous):

integral of sec^3(theta)

OpenStudy (anonymous):

any hint? Im guessing u have 2 multiply top/bottom by something

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integral+of+secx%5E3dx its a special reduction formula....

OpenStudy (anonymous):

right.. but how u go about doing it ?

OpenStudy (turingtest):

you don't need a reduction formula for this one, but the other way is a bit annoying

OpenStudy (anonymous):

got, reasoned it out. thanks

OpenStudy (turingtest):

\[\sec^3\theta=\sec^2\theta\cdot\sec\theta=(\tan^2+1)\sec\theta\]ok, if you got it already then congratulations

OpenStudy (anonymous):

ye, i did it that way :P ty

OpenStudy (turingtest):

or you could use integration by parts: http://tutorial.math.lamar.edu/Classes/CalcII/IntegralsWithTrig.aspx example 10^

OpenStudy (anonymous):

i meant inte. by part.

OpenStudy (turingtest):

that is the better way, yeah I'm just tired...

OpenStudy (anonymous):

haahahaha ya, i kno the feeling

OpenStudy (anonymous):

i got a curious question. For polar equations, When they ask u to find the point where it's horiz/verti tangents, y do they only wants the first few thetas?

OpenStudy (turingtest):

by "the first few theta's" you mean in the interval [0,2pi] ?

OpenStudy (anonymous):

no, they dont give u interval, just r=something, and that's it, when i solved it, for ex. cos(theta)=1, so theta has many value, but they always pick the first 2-3 theta (and positive)

OpenStudy (anonymous):

so i kept having extra answer, which im not sure even if it's right or wat.

OpenStudy (turingtest):

they are choosing what are called the "principle values" of theta, which are the values that satisfy in the interval \([0,2\pi]\) you can always add \(2\pi k| k\in\mathbb N\) but that goes for all values of theta in ases like this, so we ignore that part often and just write the "principle values" i.e. those between [0,2pi]

OpenStudy (turingtest):

in cases*

OpenStudy (anonymous):

so those extra one i have r ok? or should i remove them?

OpenStudy (anonymous):

wait.. after 2pi, it reapeats.. is that y ? u ignore the rest since it gona be same asnwer with different theta value?

OpenStudy (turingtest):

for the most part, yes if you had \(\frac\theta2\) as an argument it may be sightly different depends on what they are... this integral is indefinite, so it doesn't matter but \(\cos(\frac\pi2-\theta)=\sin\theta\) so you can use that identity as well in short, there are infinitely many answers, because we can go around the unit circle infinitely many times and get the same value for \(\cos\theta\), so we often just write the "principle values" i.e. those in the interval \([0,2\pi]\)

OpenStudy (turingtest):

\[\cos(\theta)=\cos(\theta+2\pi k);k\in\mathbb N\]but we exclude the second part often, because all it says is "go all the way around the unit circle and the trig function is the same" which is obvious

OpenStudy (anonymous):

got cha, I still need to memorize the unit circle, which maybe a lot of help. well ty alot.

OpenStudy (turingtest):

no prob :)